Question 6-

An entrepreneur founded his company using $300,000 of his own​ money, issuing himself 300,000 shares of stock. An angel investor bought an additional 100,000 shares for $250,000. The entrepreneur now sells another 450,000 shares of stock to a venture capitalist for $1.0 million. What is the​ post-money valuation of the​ company?

A. $2,172,222

B. $944,444

C. $1,888,889

D. $1,038,889
Answer #1

C. $1,888,889

Value of 450,000 shares is $ 1 million.

Value of 1 share is = 1 million / 450,000 = $2.2222222

Total number of shares = 450,000 + 300,000 + 100,000 =850,000

Total value of shares = 2.2222222 * 850,000 =$1,888,889

Post money valuations of company is 1,888,889 dollars